# Medverse: A Universal Model for Full-Resolution 3D Medical Image Segmentation, Transformation and Enhancement

### v1

@@ Abstract
In-context learning (ICL) offers a promising paradigm for universal medical image analysis, enabling models to perform diverse image processing tasks without retraining. However, current ICL models for medical imaging remain limited in two critical aspects: they cannot simultaneously achieve high-fidelity predictions and global anatomical understanding, and there is no unified model trained across diverse medical imaging tasks (e.g., segmentation and enhancement) and anatomical regions. As a result, the full potential of ICL in medical imaging remains underexplored. Thus, we present Medverse, a universal ICL model for 3D medical imaging, trained on 22 datasets covering diverse tasks in universal image segmentation, transformation, and enhancement across multiple organs, imaging modalities, and clinical centers. Medverse employs a next-scale autoregressive in-context learning framework that progressively refines predictions from coarse to fine, generating consistent, full-resolution volumetric outputs and enabling multi-scale anatomical awareness. We further propose a blockwise cross-attention module that facilitates long-range interactions between context and target inputs while preserving computational efficiency through spatial sparsity. Medverse is extensively evaluated on a broad collection of held-out datasets covering previously unseen clinical centers, organs, species, and imaging modalities. Results demonstrate that Medverse substantially outperforms existing ICL baselines and establishes a novel paradigm for in-context learning. Code and model weights will be made publicly available. Our model are publicly available at https://github.com/jiesihu/Medverse.
@@ Introduction
Universal medical imaging models trained on large-scale, diverse datasets represent a significant step toward clinical AI deployment, exhibiting strong generalization across clinical centers and medical tasks (Liu et al. 2023; Butoi et al. 2023; Yang et al. 2024).
Initially introduced in natural language processing (Brown et al. 2020), in-context learning (ICL) has recently emerged as a unified paradigm in medical imaging (Butoi et al. 2023; Gao et al. 2025). It conditions on a set of image–label pairs from other subjects, known as the context, to convey task information. By varying these context examples, a single model can perform diverse tasks such as segmentation, transformation, and enhancement, even including tasks that were not encountered during training (Czolbe and Dalca 2023). Unlike traditional adaptation approaches, ICL adapts to distribution shifts and novel tasks without retraining (Chan et al. 2022; Reddy 2023), making it particularly suited for heterogeneous medical imaging scenarios.
@@
SegGPT (Wang et al. 2023b) and Painter (Wang et al. 2023a) have demonstrated the effectiveness of ICL in natural image domains. However, their performance is suboptimal when applied to medical images (Hu et al. 2025). Models such as UniverSeg (Butoi et al. 2023) and One-Prompt (Wu and Xu 2024), trained on medical imaging data, have shown promising results on 2D medical image segmentation tasks. Neuralizer (Czolbe and Dalca 2023) further extends ICL to a wider range of 2D tasks, including denoising and skull stripping. However, when applied to 3D data, they suffer from the loss of topological correlations and volumetric context. More recently, Neuroverse3D (Hu et al. 2025) has extended ICL to 3D neuroimaging by introducing adaptive context processing and architectural designs addressing the computational challenges of volumetric data.
@@
Unlike conventional ICL models that utilize only context from other subjects, our NA-ICL framework provides the model with both image-label pairs from other subjects and the model’s coarse predictions at lower resolutions, inspired by (Tian et al. 2024). This architectural design enables progressive processing from coarse to fine scales, offering benefits including the integration of multi-scale anatomical information, improved consistency across patches, and high-fidelity outputs.
@@
where $F(\cdot)$ denotes the prediction function of Medverse. To efficiently handle a large number of semantic context examples, we incorporate the adaptive parallel-sequential context processing module introduced in (Hu et al. 2025), which enables scalable context processing with minimal memory requirements.
@@
The computational complexity is thereby reduced to $\mathcal{O}(B^{2})$, in contrast to the original $\mathcal{O}\!\bigl{(}(H^{\prime}W^{\prime}D^{\prime})^{2}\bigr{)}$, while still preserving long-range interactions and retaining fine-grained value information. $B=p^{3}$ is a user-defined constant that remains fixed across U-Net stages, further ensuring stable computational cost. Compared to the direct concatenation-based fusion used in (Butoi et al. 2023; Czolbe and Dalca 2023; Hu et al. 2025), this cross-attention fusion approach is better suited for handling spatial misalignment between the context and target objects.
@@
For fair comparison, we adopt the same loss functions as those proposed in (Hu et al. 2025). Specifically, a modified $\mathcal{L}_{\text{smooth}-L_{1}}$ loss is used for segmentation tasks, while for image enhancement and transformation tasks, the $\mathcal{L}_{\text{smooth}-L_{1}}$ loss is applied to both image intensity and intensity differences.
@@ Methods
Table 1: Segmentation comparison across unseen centers, organs, species, and modalities in terms of Dice scores (%). The context set and the training set for few-shot models are both of size 4.
@@
Table 2: Performance comparison across image transformation and enhancement tasks with a context size of 4 in terms of PSNR. Enhancement tasks are the average over 5 held-out datasets.
@@
Datasets. To ensure robust cross-center generalization and data diversity, we curated a collection of 27 publicly available datasets spanning multiple imaging modalities and acquisition centers, comprising a total of 40,362 3D scans. The dataset encompasses widely used medical imaging modalities, including T1, T2, FLAIR, MRA, DWI, ADC, PD, and CT, as well as commonly studied anatomical regions such as the brain, abdomen, prostate, and lung. 22 datasets (Yang et al. 2023; CAS 2023; Hernandez Petzsche et al. 2022; Liew et al. 2022; IXI 2015; Hoopes et al. 2022; Marcus et al. 2007; Flanders et al. 2020; ADHD 2011; Jack Jr et al. 2008; Alexander et al. 2017; Holmes et al. 2015; Gera et al. 2023; Nugent et al. 2022; Sudlow et al. 2015; Menze et al. 2014; Litjens et al. 2014; Kuijf et al. 2019; Zeng et al. 2023; Ji et al. 2022; Luo et al. 2024; Wasserthal et al. 2023; Antonelli et al. 2022) were used for training and validation with a random 9:1 split. The remaining five datasets were reserved as held-out sets to evaluate generalization to unseen distributions. These held-out datasets include brain and abdomen scans from previously unseen centers (Marek et al. 2011; Ma et al. 2024), the nasal cavity as an unseen anatomical target (Zhang et al. 2024), mice as an unseen species (Rosenhain et al. 2018), and PET as an unseen imaging modality (Jack Jr et al. 2008). Each held-out dataset was split in a 5:5 ratio into a meta context set used for semantic context selection and a test set.
@@
Medverse. Both the target and context branches are implemented using a five-stage 3D U-Net architecture (Çiçek et al. 2016). Each branch begins with 32 channels in the first stage, with the number of channels doubling at each subsequent stage. The model operates on input patches of size $128\times 128\times 128$. In BAM, $p$ is set to 4, and $m$ is set to 66.
Task-Specific Models. We compared Medverse against task-specific 3D models. Each task-specific model was trained directly on the meta context set of the held-out datasets, thereby mitigating domain shift concerns. This setting reflects the conventional practice in clinical environments, where models are trained specifically for a given task. Two backbone architectures were evaluated: a 3D U-Net, which shares the same architecture and channel configuration as the Medverse backbone, and the Swin-UNETR (Hatamizadeh et al. 2021). We assessed performance under both few-shot and fully supervised learning scenarios.
Other ICL Models. We compared our method with several state-of-the-art ICL approaches, including Painter (Wang et al. 2023a), SegGPT (Wang et al. 2023b), UniverSeg (Butoi et al. 2023), and Neuralizer (Czolbe and Dalca 2023), all of which are designed for 2D inputs. To adapt to 3D, we split each 3D target image into 2D slices and randomly sampled slices from the 3D context set containing the target region to construct the 2D context set. The number of context slices was set to 1 for Painter, 8 for SegGPT, 32 for Neuralizer, and 64 for UniverSeg, following the optimal configurations reported in their respective publications. The resulting 2D outputs were reassembled into 3D volumes for metric evaluation. We also compared with Neuroverse3D (Hu et al. 2025), an ICL model developed for 3D neuroimaging. Due to its limited input resolution of $128\times 128\times 128$, we resized the input images to fit Neuroverse3D and rescaled the outputs back to the original resolution for evaluation. All models were evaluated using their publicly released pretrained weights.
@@
Ablation Analysis of Key Components. Table 3 shows ablations on the BAM module and the NA-ICL framework. Removing NA-ICL, which replaces coarse-to-fine prediction with direct sliding-window inference, leads to a notable drop in segmentation performance. This is due to the limited perspective field of the target image and the loss of global context guidance. For transformation and enhancement tasks, NA-ICL also improves performance, with one contributing factor being its ability to enhance consistency across sliding-window patches. When the BAM module is removed, feature fusion is instead performed via simple concatenation as in (Hu et al. 2025). The BAM module leads to overall performance improvements due to mitigating the adverse effects caused by spatial misalignment between the context and target images.

### v3

@@ Abstract
In-context learning (ICL) offers a promising paradigm for universal medical image analysis, enabling models to perform diverse image processing tasks without retraining. However, current ICL models for medical imaging remain limited in two critical aspects: they cannot simultaneously achieve high-fidelity predictions and global anatomical understanding, and there is no unified model trained across diverse medical imaging tasks (e.g., segmentation and enhancement) and anatomical regions. As a result, the full potential of ICL in medical imaging remains underexplored. Thus, we present Medverse, a universal ICL model for 3D medical imaging, trained on 22 datasets covering diverse tasks in universal image segmentation, transformation, and enhancement across multiple organs, imaging modalities, and clinical centers. Medverse employs a next-scale autoregressive in-context learning framework that progressively refines predictions from coarse to fine, generating consistent, full-resolution volumetric outputs and enabling multi-scale anatomical awareness. We further propose a blockwise cross-attention module that facilitates long-range interactions between context and target inputs while preserving computational efficiency through spatial sparsity. Medverse is extensively evaluated on a broad collection of held-out datasets covering previously unseen clinical centers, organs, species, and imaging modalities. Results demonstrate that Medverse substantially outperforms existing ICL baselines and establishes a novel paradigm for in-context learning. Our model are publicly available at https://github.com/jiesihu/Medverse.
@@ Introduction
Universal medical imaging models trained on large-scale, diverse datasets represent a significant step toward clinical AI deployment, exhibiting strong generalization across clinical centers and medical tasks (liu2023clip; butoi2023universeg; yang2024brainmass).
Initially introduced in natural language processing (brown2020language), in-context learning (ICL) has recently emerged as a unified paradigm in medical imaging (butoi2023universeg; gao2025show). It conditions on a set of image–label pairs from other subjects, known as the context, to convey task information. By varying these context examples, a single model can perform diverse tasks such as segmentation, transformation, and enhancement, even including tasks that were not encountered during training (czolbe2023neuralizer). Unlike traditional adaptation approaches, ICL adapts to distribution shifts and novel tasks without retraining (chan2022data; reddy2023mechanistic), making it particularly suited for heterogeneous medical imaging scenarios.
@@
SegGPT (wang2023seggpt) and Painter (wang2023images) have demonstrated the effectiveness of ICL in natural image domains. However, their performance is suboptimal when applied to medical images (hu2025building). Models such as UniverSeg (butoi2023universeg) and One-Prompt (wu2024one), trained on medical imaging data, have shown promising results on 2D medical image segmentation tasks. Neuralizer (czolbe2023neuralizer) further extends ICL to a wider range of 2D tasks, including denoising and skull stripping. However, when applied to 3D data, they suffer from the loss of topological correlations and volumetric context. More recently, Neuroverse3D (hu2025building) has extended ICL to 3D neuroimaging by introducing adaptive context processing and architectural designs addressing the computational challenges of volumetric data.
@@
Unlike conventional ICL models that utilize only context from other subjects, our NA-ICL framework provides the model with both image-label pairs from other subjects and the model’s coarse predictions at lower resolutions, inspired by (tian2024visual). This architectural design enables progressive processing from coarse to fine scales, offering benefits including the integration of multi-scale anatomical information, improved consistency across patches, and high-fidelity outputs.
@@
where $F(\cdot)$ denotes the prediction function of Medverse. To efficiently handle a large number of semantic context examples, we incorporate the adaptive parallel-sequential context processing module introduced in (hu2025building), which enables scalable context processing with minimal memory requirements.
@@
The computational complexity is thereby reduced to $\mathcal{O}(B^{2})$, in contrast to the original $\mathcal{O}\!\bigl((H^{\prime}W^{\prime}D^{\prime})^{2}\bigr)$, while still preserving long-range interactions and retaining fine-grained value information. $B=p^{3}$ is a user-defined constant that remains fixed across U-Net stages, further ensuring stable computational cost. Compared to the direct concatenation-based fusion used in (butoi2023universeg; czolbe2023neuralizer; hu2025building), this cross-attention fusion approach is better suited for handling spatial misalignment between the context and target objects.
@@
For fair comparison, we adopt the same loss functions as those proposed in (hu2025building). Specifically, a modified $\mathcal{L}_{\text{smooth}-L_{1}}$ loss is used for segmentation tasks, while for image enhancement and transformation tasks, the $\mathcal{L}_{\text{smooth}-L_{1}}$ loss is applied to both image intensity and intensity differences.
@@ Methods
Table 1: Segmentation comparison across unseen centers, organs, species, and modalities in terms of Dice scores (%). The context set and the training set for few-shot models are both of size 4. 2D models marked with an * utilize the orthogonal view fusion method.
@@
Table 2: Performance comparison across image transformation and enhancement tasks with a context size of 4 in terms of PSNR. Enhancement tasks are the average over 5 held-out datasets. 2D models marked with an * utilize the orthogonal view fusion method.
@@
Datasets. To ensure robust cross-center generalization and data diversity, we curated a collection of 27 publicly available datasets spanning multiple imaging modalities and acquisition centers, comprising a total of 40,362 3D scans. The dataset encompasses widely used medical imaging modalities, including T1, T2, FLAIR, MRA, DWI, ADC, PD, and CT, as well as commonly studied anatomical regions such as the brain, abdomen, prostate, and lung. 22 datasets (yang2023benchmarking; CAS2023; hernandez2022isles; liew2022large; IXI; hoopes2022learning; marcus2007open; flanders2020construction; ADHD; jack2008alzheimer; alexander2017open; holmes2015brain; gera2023characterizing; nugent2022nimh; sudlow2015uk; menze2014multimodal; litjens2014evaluation; kuijf2019standardized; zeng2023imagecas; ji2022amos; luo2024rethinking; wasserthal2023totalsegmentator; antonelli2022medical) were used for training and validation with a random 9:1 split. The remaining five datasets were reserved as held-out sets to evaluate generalization to unseen distributions. These held-out datasets include brain and abdomen scans from previously unseen centers (marek2011parkinson; ma2024unleashing), the nasal cavity as an unseen anatomical target (zhang2024nasalseg), mice as an unseen species (rosenhain2018preclinical), and PET as an unseen imaging modality (jack2008alzheimer). Each held-out dataset was split in a 5:5 ratio into a meta context set used for semantic context selection and a test set.
@@
Medverse. Both the target and context branches are implemented using a five-stage 3D U-Net architecture (cciccek20163d). Each branch begins with 32 channels in the first stage, with the number of channels doubling at each subsequent stage. The model operates on input patches of size $128\times 128\times 128$. In BAM, $p$ is set to 4, and $m$ is set to 66.
Task-Specific Models. We compared Medverse against task-specific 3D models. Each task-specific model was trained directly on the meta context set of the held-out datasets, thereby mitigating domain shift concerns. This setting reflects the conventional practice in clinical environments, where models are trained specifically for a given task. Two backbone architectures were evaluated: a 3D U-Net, which shares the same architecture and channel configuration as the Medverse backbone, and the Swin-UNETR (hatamizadeh2021swin). We assessed performance under both few-shot and fully supervised learning scenarios.
Other ICL Models. We compared our method with several state-of-the-art ICL approaches, including Painter (wang2023images), SegGPT (wang2023seggpt), UniverSeg (butoi2023universeg), and Neuralizer (czolbe2023neuralizer), all of which are designed for 2D inputs. To adapt to 3D, we split each 3D target image into 2D slices and randomly sampled slices from the 3D context set containing the target region to construct the 2D context set. The number of context slices was set to 1 for Painter, 8 for SegGPT, 32 for Neuralizer, and 64 for UniverSeg, following the optimal configurations reported in their respective publications. The resulting 2D outputs were reassembled into 3D volumes for metric evaluation. 2D models marked with an asterisk (*) utilize the orthogonal view fusion method from 3 views, which introduces a three-fold increase in computational load. We also compared with Neuroverse3D (hu2025building), an ICL model developed for 3D neuroimaging. Due to its limited input resolution of $128\times 128\times 128$, we resized the input images to fit Neuroverse3D and rescaled the outputs back to the original resolution for evaluation. All models were evaluated using their publicly released pretrained weights.
@@
Ablation Analysis of Key Components. Table 3 shows ablations on the BAM module and the NA-ICL framework. Removing NA-ICL, which replaces coarse-to-fine prediction with direct sliding-window inference, leads to a notable drop in segmentation performance. This is due to the limited perspective field of the target image and the loss of global context guidance. For transformation and enhancement tasks, NA-ICL also improves performance, with one contributing factor being its ability to enhance consistency across sliding-window patches. When the BAM module is removed, feature fusion is instead performed via simple concatenation as in (hu2025building). The BAM module leads to overall performance improvements due to mitigating the adverse effects caused by spatial misalignment between the context and target images.